Ahlheim Corporation has two production departments, Forming and Assembly. The company uses a job-order costing system and computes a predetermined overhead rate in each production department. The Forming Department’s predetermined overhead rate is based on machine-hours and the Assembly Department’s predetermined overhead rate is based on direct labor-hours. At the beginning of the current year, the company had made the following estimates:

Forming Assembly
Machine-hours 16,000 15,000
Direct labor-hours 2,000 6,000
Total fixed manufacturing overhead cost $ 102,400 $ 55,200
Variable manufacturing overhead per machine-hour $ 2.30
Variable manufacturing overhead per direct labor-hour $ 4.50

During the current month the company started and finished Job T924. The following data were recorded for this job:

Job T924: Forming Assembly
Machine-hours 70 20
Direct labor-hours 30 40
Direct materials $ 870 $ 385
Direct labor cost $ 630 $ 840

The estimated total manufacturing overhead for the Assembly Department is closest to:

Multiple Choice

  • $27,000

  • $82,200

  • $55,200

  • $47,700
Answer #1

Answer : $ 82,200

Estimated Manufacturing overhead = Fixed Manufacturing overhead + ( overhead *Direct labor hours)

Estimated Manufacturing overhead

= 55,200+(4.50*6000)= 55,200+27,000 =$ 82,200 (Answer)